OpenStudy (anonymous):

x^2-x-2/2x^2-5x+2 (divided by) x^2-x-12/2x^2+5x-3

zepdrix (zepdrix):

wow this problem is evil :D algebra nightmare lol hmm ok lets see..

zepdrix (zepdrix):

remember that when you divide fractions, it's the same as multiplying by the reciprocal of the denominator. you take the bottom fraction and "flip it" and rewrite it as multiplication

OpenStudy (anonymous):

im more confused :-( @zepdrix

zepdrix (zepdrix):

so you can rewrite this equation like so \[\frac{ x^{2}-x-2 }{ 2x^{2}-5x+2 }*\frac{ 2x^{2}+5x-3 }{ x^{2}-x-12 }\]

zepdrix (zepdrix):

D':

zepdrix (zepdrix):

\[\frac{ (\frac{ a }{ b }) }{ (\frac{ c }{ d }) }\] --->\[\frac{ a }{ b }*\frac{ d }{ c }\]

zepdrix (zepdrix):

see what i did there? <:O the bottom fraction flips and the problem becomes multiplication instead of dividing 2 fractions

OpenStudy (anonymous):

:-( @zepdrix

zepdrix (zepdrix):

lol :D

OpenStudy (anonymous):

im still confused can u do it step by step @zepdrix

zepdrix (zepdrix):

hmm imma draw picture, its too hard to keep typing out that thing lol

zepdrix (zepdrix):

 ok this is the first step D: so try to wrap yer head around it. i rewrote the example using simple letters next to it